# CHIMERA: A Knowledge Base of Idea Recombination in Scientific Literature

### v3

@@ Appendix C E2E vs Specialized extraction
This section reflects on the results described in Section 4, drawing on implementation details of the baselines (described in Appendix B). In Section 4, we observe that narrowing the focus to a smaller portion of the recombination extraction task does not always improve performance - in fact, it can lead to worse results. This pattern emerges across three Mistral-based classifiers: the end-to-end version (E2E), the specialized version (Abstract-classifier), and the specialized version trained with synthetic CoT strings (Abstract-classifier-CoT). We hypothesize that identifying recombination relations in text may be analogous to Chain-of-Thought prompting (CoT), a technique known to enhance LLM performance across various tasks Wei et al. (2022). This hypothesis is supported by the superior performance of Abstract-classifier-CoT compared to its non-CoT counterpart.
Appendix D Graph nodes domains
@@
Figure 10: blend domain analysis prompt. {ELEMENTS} is a placeholder for the recombination entities extracted from {ABSTRACT}. {ARXIV} is a placeholder for full arXiv category names and their descriptions. {BRANCHES} is a placeholder for the list of non-arXiv domains given in Appendix D, Table 8.
Figure 11: inspiration domain analysis prompt. {INSPIRATION_SOURCE} and {INSPIRATION_TARGET} are placeholders for the inspiration entities extracted from {ABSTRACT}. {ARXIV} is a placeholder for full arXiv category names and their descriptions. {BRANCHES} is a placeholder for the list of non-arXiv domains given in Appendix D, Table 8.
We identify entities’ scientific domain using GPT-4o in zero-shot settings. Given the abstract and recombination entities extracted from it, the model has to assign each an arXiv category and a scientific branch. In case the model manages to assign the entity an arxive category, the scientific branch is the category’s full name (e.g., "Artificial Intelligence" for cs.AI). Otherwise, the models assign the branch from a list of outer-arXiv domains described in Table 8. If the model can assign the entity a standard arXiv category, we use it as the domain. Otherwise, we use the branch (an outer-arXiv domain). Entities with neither are assigned to the Other domain. Figures 10 and 11 present our analysis prompts for blend and inspiration relations, respectively. The cost of running the analysis over the collected corpus was 250\$.
@@
We use the Other domain for nodes the model fails to analyze, and $2127$ of the graph nodes are assigned to this category. We examine a sample of 150 such nodes and observe that many are too noisy or overly general to classify. Interestingly, some nodes in this domain describe nonacademic or niche concepts (see examples in Table 9).
@@
Appendix E Span similarity prompt
We provide our span similarity prompt in Figure E. We use it in the extraction evaluation process as discussed in Section 4.1.
Appendix F Error analysis
Bad extraction examples
Pred = []
[…]
Pred = []
Context dependent or implicit phrasing
We observe that, unsurprisingly, cases in which the recombination is implied or subtle are more challenging for the model. For instance (see also Table 12, row 1), "Kahneman & Tversky’s prospect theory" inspires the design of a loss function that "directly maximizes the utility of generations", but this is not stated explicitly. Moreover, abstracts that explicitly express idea recombination while referencing previously mentioned entities are also harder to detect.
Multiple recombinations
Borderline cases
Appendix G Extraction examples
Detected Inspiration Abstract
Inspiration-Source:
"the shepherding behavior of herding dogs" [zoology]
Inspiration-Target:
"Frontier exploration" [cs.ro]
Inspiration-Source:
Inspiration-Target:
"Histopathological image classification" [cs.cv]
Inspiration-Source:
"the dual-process theory in psychology" [psychology]
Inspiration-Target:
"enhance dialogue planning in LLMs" [cs.cl]
Inspiration-Source:
Inspiration-Target:
"learning effective feature embeddings for CTR prediction" [cs.lg]
Appendix H Predominant inspiration and blend relations
@@
Appendix I Prediction data preprocessing
@@
We use GPT-4o-mini to extract a few sentences from each abstract describing the background or motivation of the authors using recombination (See prompt on Figure 13). Adding these contexts to the queries helps them be more specific and limits the search space. However, this might introduce leaks into the queries - cases where the extracted context reveals the answer. Table 15 presents leak examples. We utilize GPT-4o-mini again to filter out such cases from the data, using the prompt shown in Figure 14. In a qualitative analysis of 50 randomly sampled query-answer pairs, we find that a human annotator agrees with 87% of the model’s predictions (whether there is a leak). Finally, we divide the remaining query-answer pairs into splits as described in Table 5 is Section 5.2.
Appendix J Prediction baselines
@@
We use a bi-encoder architecture for recombination prediction and experiment with three popular encoders as backbones: all-mpnet-base-v2 (109M parameters), bge-large-en-v1.5 Xiao et al. (2023) (335M parameters) and e5-large-v2 Wang et al. (2022) (335M parameters). These models’ checkpoints predate 2024, meaning they are unfamiliar with our test set. The model receives a query string composed of a context description, a graph entity, and a relation type and returns a ranked list of answers (other graph nodes). We perform HPO (random grid search of 10 trails) to select the number of training epochs, warmup ratio and learning rate for each model. We use contrastive loss and generate 30 negatives per positive example. Following the literature standard Teach et al. (2020), we report metrics in the filtered settings to avoid false negatives. Given the difficulty of the task we focus on ranking only the $12751$ test set entities. A full summary of our data splits is available on 5. The examples we use to train and evaluate our prediction models contain all collected nodes, including those classified as belonging to the "other" domain.
@@
Appendix K User study additional details
@@
Appendix L Comparison to other information extraction methods

### v4

@@ B.3 E2E vs Specialized extraction
This section reflects on the results described in Section 4, drawing on implementation details of the baselines (described in Appendix B.2). In Section 4, we observe that narrowing the focus to a smaller portion of the recombination extraction task does not always improve performance - in fact, it can lead to worse results. This pattern emerges across three Mistral-based classifiers: the end-to-end version (E2E), the specialized version (Abstract-classifier), and the specialized version trained with synthetic CoT strings (Abstract-classifier-CoT). We hypothesize that identifying recombination relations in text may be analogous to Chain-of-Thought prompting (CoT), a technique known to enhance LLM performance across various tasks Wei et al. (2022). This hypothesis is supported by the superior performance of Abstract-classifier-CoT compared to its non-CoT counterpart.
B.4 Span similarity
We provide our span similarity prompt in Figure B.4. We use it in the extraction evaluation process as discussed in Section 4.1. To mitigate position bias, we query the model twice per pair with reversed orderings, accepting a match only if both judgments are positive. We prefer GPT-4o-mini over GPT-4o based on a comparison which found only 333 disagreements across the test set.
B.5 Extraction error analysis
Bad extraction examples (human annotated test set)
Pred = []
[…]
Pred = []
Context dependent or subtle phrasing
We observe that, unsurprisingly, cases in which the recombination is implied or subtle are more challenging for the model. For instance (see also Table 9, row 1), "Kahneman & Tversky’s prospect theory" inspires the design of a loss function that "directly maximizes the utility of generations", but this is not stated directly. Moreover, abstracts that express idea recombination while referencing previously mentioned entities are also harder to detect.
Multiple recombinations
Borderline cases
B.6 Extraction examples
Detected Inspiration Abstract
Inspiration-Source:
"the shepherding behavior of herding dogs" [zoology]
Inspiration-Target:
"Frontier exploration" [cs.ro]
Inspiration-Source:
Inspiration-Target:
"Histopathological image classification" [cs.cv]
Inspiration-Source:
"the dual-process theory in psychology" [psychology]
Inspiration-Target:
"enhance dialogue planning in LLMs" [cs.cl]
Inspiration-Source:
Inspiration-Target:
"learning effective feature embeddings for CTR prediction" [cs.lg]
Bad Extraction Examples (arXiv)
Abstract: "Supervised deep learning with pixel-wise training labels has great successes on multi-person part segmentation. However, data labeling at pixel-level is very expensive. To solve the problem, people have been exploring to use synthetic data…the results are much worse compared to those using real data and manual labeling. The degradation of the performance is mainly due to the domain gap, i.e., the discrepancy of the pixel value statistics between real and synthetic data. In this paper, we observe that real and synthetic humans both have a skeleton (pose) representation. We found that the skeletons can effectively bridge the synthetic and real domains during the training. Our proposed approach takes advantage of the rich and realistic variations of the real data and the easily obtainable labels of the synthetic data to learn multi-person part segmentation on real images without any human-annotated labels… "
[…]
Abstract: "…In this paper we propose an LLM feature-based framework for dialogue constructiveness assessment that combines the strengths of feature-based and neural approaches, while mitigating their downsides. The framework first defines a set of dataset-independent and interpretable linguistic features, which can be extracted by both prompting an LLM and simple heuristics. Such features are then used to train LLM feature-based models…We also find that the LLM feature-based model learns more robust prediction rules instead of relying on superficial shortcuts, which often trouble neural models."
[…]
Table 11: Representative examples of bad automatic extraction. Many errors stem from uninformative entity spans, as presented by the two bottom examples.
B.7 Large-scale extraction assessment
Figure 12: Large-scale evaluation prompt. {ABSTRACT} is a placeholder for the original abstract text. {EXTRACTED_RELATION}, {ENTITY1}, and {ENTITY2} are placeholders for the relation type and entities extracted by our model.
To complement our human annotation efforts and enable large-scale evaluation, we conducted a qualitative assessment of the automatically extracted recombination examples in CHIMERA using GPT-4.1 as an LLM-based judge.
Validating the LLM Judge.
We first assessed GPT-4.1’s reliability by comparing its judgments against those of a domain expert. A PhD student with NLP expertise manually reviewed 100100100 randomly sampled recombination examples and labeled each as correct if: (1) the extracted entities corresponded to meaningful scientific concepts, and (2) the relation between them captured a central recombination explicitly described in the abstract. Upon analyzing the identified extraction errors, we observe a significant portion stems from extracting correct recombinations with uninformative entities (criteria 2) and not from a conceptual misunderstanding of the text. We provide examples of such cases in Table B.6.
GPT-4.1 was prompted with the same examples using an evaluation template aligned with the assessment criteria (see Figure 12). Given the imbalance nature of the data, we report the F1 score instead of Cohen’s $\kappa$, following the recommendations of previous work Delgado and Tibau (2019). The resulting F1 score of 0.9120.9120.912 indicates substantial agreement, supporting the use of GPT-4.1 as a reliable proxy for large-scale quality assessment.
Large-Scale Evaluation.
Following validation, we applied GPT-4.1 to a larger sample of 2,000 automatically extracted examples from CHIMERA. The model labeled 799 of these examples as correct, resulting in an estimated extraction accuracy of 80.55%. These results provide further evidence for the overall quality and robustness of our extraction pipeline.
Appendix C Graph nodes domains
@@
Figure 13: blend domain analysis prompt. {ELEMENTS} is a placeholder for the recombination entities extracted from {ABSTRACT}. {ARXIV} is a placeholder for full arXiv category names and their descriptions. {BRANCHES} is a placeholder for the list of non-arXiv domains given in Appendix C, Table 12.
Figure 14: inspiration domain analysis prompt. {INSPIRATION_SOURCE} and {INSPIRATION_TARGET} are placeholders for the inspiration entities extracted from {ABSTRACT}. {ARXIV} is a placeholder for full arXiv category names and their descriptions. {BRANCHES} is a placeholder for the list of non-arXiv domains given in Appendix C, Table 12.
We identify the scientific domain of each entity using GPT-4o in a zero-shot setting. Given the abstract and the extracted recombination entities, the model assigns to each entity an arXiv category and a broader scientific branch. If the model successfully assigns an arXiv category, we treat it as the entity’s domain.
Otherwise, the model selects a branch from a predefined list of outer-arXiv domains (see Table 12) and sets it as the domain. If neither a standard arXiv category nor a branch can be assigned, the entity is labeled as belonging to the Other domain.
Entities in the Other domain are excluded from the analysis in Section 4.3, as they are often too noisy, overly broad, or miscellaneous to interpret reliably. Figures 13 and 14 present the prompts used for analyzing blend and inspiration relations, respectively. Running this classification process over the full corpus cost approximately 250\$.
@@
We assign the Other domain to nodes the model fails to classify. In total, 2,127 graph nodes fall into this category. We manually examined a sample of 150 such nodes and found that many were either too ambiguous or too general to categorize meaningfully. Interestingly, some of these nodes refer to non-academic or highly niche concepts (see examples in Table 13).
@@
Appendix D Additional Knowledge Base Analysis
D.1 Predominant recombination relations
@@
D.2 Nuanced recombination types
Nuanced recombination types examples
Reduction
Abstract: "Register allocation is one of the most important problems for modern compilers…This work demonstrates the use of casting the register allocation problem as a graph coloring problem…"
Inspiration: "a graph coloring problem" $\longrightarrow$ "Register allocation"
Metaphor
Abstract: "Affective sharing within groups strengthens coordination and empathy…we propose HeartBees, a bio-feedback system for visualizing collective emotional states, which maps a multi-dimensional emotion model into a metaphorical visualization of flocks of birds…"
Inspiration: "flocks of birds" $\longrightarrow$ "a multi-dimensional emotion model"
Analogy
Abstract: "Physics-informed Graph Neural Networks have achieved remarkable performance…by mitigating common GNN challenges…Despite these advancements, the development of a simple yet effective paradigm that appropriately integrates previous methods for handling all these challenges is still underway. In this paper, we draw an analogy between the propagation of GNNs and particle systems in physics, proposing a model-agnostic enhancement framework…"
Inspiration: "particle systems in physics" $\longrightarrow$ "the propagation of GNNs"
Application of existing knowledge
Abstract: "Object detection in high-resolution satellite imagery is emerging as a scalable alternative to on-the-ground survey data collection…However, performing object detection over large geographies can still be prohibitively expensive due to the high cost of purchasing imagery and compute. Inspired by traditional survey data collection strategies, we propose an approach to estimate object count statistics over large geographies through sampling…"
Inspiration: "traditional survey data collection strategies" $\longrightarrow$ "Object detection in high-resolution satellite imagery"
Abstraction
Abstract: "While visual question-answering (VQA) benchmarks have catalyzed the development of reasoning techniques, they have focused on vertical thinking. Effective problem-solving also necessitates lateral thinking…To bridge this gap, we formulate visual lateral thinking as a multiple-choice question-answering task…"
Inspiration: "a multiple-choice question-answering task" $\longrightarrow$ "visual lateral thinking"
Table 17: Examples of nuanced inspiration types found within CHIMERA. While all examples are labeled as inspiration, they illustrate finer-grained mechanisms such as metaphor, reduction, and analogy. This suggests that our taxonomy is expressive enough to capture a rich diversity of recombination strategies.
In Section 3, we defined the two broad recombination types used in CHIMERA: blends and inspirations. In this section, we demonstrate that this taxonomy is both robust and expressive, offering broad coverage of more nuanced recombination phenomena.
To support this, we perform a qualitative analysis of 303030 inspiration examples from the CHIMERA dataset. We identify distinct subtypes of inspiration, such as analogy, metaphor, reduction, abstraction, and application of existing knowledge. These subtypes emerge naturally within our current schema, illustrating its extensibility and broad coverage. Table 17 provides examples for each. This analysis lays the groundwork for future refinement and expansion of the taxonomy.
Appendix E Additional Prediction Details
E.1 Prediction data preprocessing
@@
We use GPT-4o-mini to extract a few sentences from each abstract describing the background or motivation of the authors using recombination (See prompt on Figure 15). Adding these contexts to the queries helps them be more specific and limits the search space. However, this might introduce leaks into the queries - cases where the extracted context reveals the answer. Table 18 presents leak examples. We utilize GPT-4o-mini again to filter out such cases from the data, using the prompt shown in Figure 16. In a qualitative analysis of 50 randomly sampled query-answer pairs, we find that a human annotator agrees with 87% of the model’s predictions (whether there is a leak). Finally, we divide the remaining query-answer pairs into splits as described in Table 5 is Section 5.
E.2 Prediction baselines
@@
We use a bi-encoder architecture for recombination prediction and experiment with three popular encoders as backbones: all-mpnet-base-v2 (109M parameters), bge-large-en-v1.5 Xiao et al. (2023) (335M parameters) and e5-large-v2 Wang et al. (2022) (335M parameters). These models’ checkpoints predate 2024, meaning they are unfamiliar with our test set. The model receives a query string composed of a context description, a graph entity, and a relation type and returns a ranked list of answers (other graph nodes). We perform HPO (random grid search of 10 trails) to select the number of training epochs, warmup ratio and learning rate for each model. We use contrastive loss and generate 30 negatives per positive example. Following the literature standard Teach et al. (2020), we report metrics in the filtered settings to avoid false negatives. Given the difficulty of the task we focus on ranking only the 127511275112751 test set entities. A full summary of our data splits is available on 5. The examples we use to train and evaluate our prediction models contain all collected nodes, including those classified as belonging to the "other" domain.
@@
E.3 Reranker error analysis
Reranking error examples
(i) Multiple plausible answers
Query: "Traditional reasoning methods in language models often rely on historical information and employ a uni-directional reasoning strategy… This leads to suboptimal decision-making… What would be a good source of inspiration for enhancing the decision rationality of language models?"
Pre-reranking (top-20)
1. the inherent human attribute of engaging in logical reasoning to facilitate decision-making
2. principles of rational decision-making
3. the Level-K framework from game theory and behavioral economics, which extends reasoning from simple reactions to structured strategic depth
Post-reranking (top-20)
1. the Level-K framework from game theory and behavioral economics, which extends reasoning from simple reactions to structured strategic depth
2. Bayesian inference: conditioning a prior on evidence
6. principles of rational decision-making
Query: "…while Large Language Models (LLMs) excel in various NLP tasks, their ability to generate comprehensive data stories remains underexplored… What would be a good source of inspiration for Data-driven storytelling?"
Pre-reranking (top-20)
1. the human storytelling process
2. story writing
3. Interactive digital stories
9. narrative structure designs
Post-reranking (top-20)
1. story analysis and generation systems
2. generative artificial intelligence (Gen-AI)-driven narrative personalization
3. narrative structure designs
4. the human storytelling process
(ii) Semantically similar variants
Query: "Prior methods for aligning large language models face challenges in tuning to maximize non-differentiable and non-binary objectives…This highlights a need for a more flexible approach that can generalize to various user preferences… while maintaining alignment… What could we blend with reinforcement learning via human feedback to address the described settings?"
Pre-reranking (top-20)
1. aligning Large Language Models with human preferences
2. Direct Preference Optimization for preference alignment
3. direct preference optimization
5. State-of-the-art language model fine-tuning techniques, such as Direct Preference Optimization
Post-reranking (top-20)
1. Direct Preference Optimization for preference alignment
2. State-of-the-art language model fine-tuning techniques, such as Direct Preference Optimization
3. contrastive learning-based methods like Direct Preference Optimization
4. a Semi-Policy Preference Optimization method
5. direct preference optimization
Table 19: Illustrative examples where the reranker preferred a different answer over the gold one.
[…]
(i) Multiple plausible answers.
In some cases, the reranker correctly identifies a strong and highly relevant candidate, and ranks it above the gold even though both answers are valid. These errors stem not from a lack of understanding, but from the presence of several equally reasonable responses. For instance, in Table 19 (top), the reranker promotes a conceptually grounded strategy from game theory over a more generic gold response about rational decision principles.
(ii) Semantically similar variants.
Another common error involves the reranker prioritizing paraphrased or reformulated versions of the gold answer. While these candidates are semantically close to the gold, the gold itself may fall in rank due to redundancy. As shown in Table 19 (bottom), several variants of "Direct Preference Optimization" receive high rankings, but the original mention of the method is pushed downward, possibly due to lexical overlap penalties or insufficient canonicalization.
These examples highlight nuanced challenges in reranking systems, such as handling redundancy and conceptual equivalence.
Appendix F User study additional details
@@
F.1 Predictions examples
Table 20 shows a selection of model predictions that participants rated as most helpful for inspiring research directions. These examples highlight how CHIMERA-trained models can move beyond surface-level associations to propose insightful cross-domain inspirations, for instance, linking harmful meme detection to visual commonsense reasoning, or drawing on neuroscience to improve LLM knowledge retention. Such predictions demonstrate CHIMERA’s potential to power ideation tools that help researchers identify novel, actionable directions for future work.
Prediction Examples
Query: Existing large language models (LLMs) underperform in legal judgment prediction due to challenges in understanding case complexities and distinguishing between similar charges. This highlights a need for improved methodologies that can effectively address these issues to enhance judicial efficiency. In this context, what would be a good source of inspiration for legal judgment prediction?
Prediction: The Issue, Rule, Application, Conclusion (Issue, Rule, Application, Conclusion) framework, a sequential reasoning approach used by lawyers.
Query: The exponential increase in scientific literature makes it challenging for researchers to stay current with recent advances and identify meaningful research directions. Existing methods for idea generation either trivially prompt large language models or directly expose them to extensive literature without indicating useful information, highlighting a need for more effective approaches to research ideation. In this context, what would be a good source of inspiration for automating the generation of novel research ideas?
Prediction: Emulate the human process of inducing collective creativity through engaging discussions with participants from diverse backgrounds and perspectives.
Query:The challenge of detecting harmful memes lies in the implicit meanings embedded within them, which are not explicitly conveyed through surface text and images. Existing harmful meme detection methods fail to provide readable explanations that reveal these implicit meanings, highlighting a significant gap in the ability to support detection decisions. In this context, what would be a good source of inspiration for detecting harmful memes?
Prediction: Visual commonsense discovery in computer vision.
Query: Large language models (LLMs) often struggle to provide up-to-date information due to their one-time training and the constantly evolving nature of the world. Existing approaches to keep LLMs current face difficulties in extracting stored knowledge, highlighting a need for improved methods of knowledge acquisition from raw documents. In this context, what would be a good source of inspiration for improving an llm’s ability to effectively acquire new knowledge from raw documents?
Prediction: Neuroscience, where the human brain often sheds outdated information to improve the retention of crucial knowledge and facilitate the acquisition of new information.
Table 20: Examples of recombination directions predicted by our model and rated as most inspiring by user study participants. Each prediction links a scientific challenge with a cross-domain concept, illustrating CHIMERA ’s potential to support creative research ideation.
Appendix G Comparison to other information extraction methods